Find the critical values x?and XR for the given confidence level c and sample size n. C=0.8, n=30 xL= (Round to three decimalFind the critical values x and x for the given confidence level c and sample size n. c=0.9, n=21 x?L= (Round to three decimalUse technology to construct the confidence intervals for the population variance o2 and the population standard deviation o.As part of a water quality survey, you test the water hardness in several randomly selected streams. The results are shown beA magazine includes a report on the energy costs per year for 32-inch liquid crystal display (LCD) televisions. The article sAs part of a water quality survey, you test the water hardness in several randomly selected streams. The results are shown be
Answer #1

df = 30 - 1 = 29

α = 1 - 0.8 = 0.2
The critical values for α = 0.2 and df = 29 are Χ^2(α/2,n-1) = 39.087


2)


n = 21
df = 21 - 1 = 20

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 20 Χ^2(α/2,n-1) = 31.410

3)

Here s = 3.7 and n = 30
df = 30 - 1 = 29

α = 1 - 0.99 = 0.01
The critical values for α = 0.01 and df = 29 are Χ^2(1-α/2,n-1) = 13.121 and Χ^2(α/2,n-1) = 52.336

phpF5A7s2.png

CI = (29*3.7^2/52.336 , 29*3.7^2/13.121)
CI = (7.59 , 30.26)

4)

Here s = 15 and n = 20
df = 20 - 1 = 19

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 19 are Χ^2(1-α/2,n-1) = 10.117 and Χ^2(α/2,n-1) = 30.144

php0mZD71.png

CI = (19*15^2/30.144 , 19*15^2/10.117)
CI = (141.8 , 422.6)

5)

Here s = 3.77 and n = 14
df = 14 - 1 = 13

α = 1 - 0.8 = 0.2
The critical values for α = 0.2 and df = 13 are Χ^2(1-α/2,n-1) = 7.042 and Χ^2(α/2,n-1) = 19.812

phpyyJ3Vy.png

CI = (13*3.77^2/19.812 , 13*3.77^2/7.042)
CI = (9.33 , 26.24)


b)

Here s = 3.77 and n = 14
df = 14 - 1 = 13

α = 1 - 0.8 = 0.2
The critical values for α = 0.2 and df = 13 are Χ^2(1-α/2,n-1) = 7.042 and Χ^2(α/2,n-1) = 19.812

CI = (sqrt(13*3.77^2/19.812) , sqrt(13*3.77^2/7.042))
CI = (3.05 , 5.12)


6)

Here s = 15 and n = 20
df = 20 - 1 = 19

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 19 are Χ^2(1-α/2,n-1) = 10.117 and Χ^2(α/2,n-1) = 30.144

php1qjxWc.png

CI = (19*15^2/30.144 , 19*15^2/10.117)
CI = (141.8 , 422.6)